# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 5.2.2 Operator Ratios From Proper Positive Semidefinite Functions
We will say that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if there exists some $k\in{\mathrm{R}}_{\nu}^{s}$ at which $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
@@
Suppose that there exists $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that $\alpha(\nu_{k})>0$ for some $k\in{\mathrm{R}}_{\nu}^{s}$ and $\alpha(T^{k}(x))\leq g(T^{k}(x))$ for all $k\in\mathbb{N}$ all $x\in X^{\mathrm{in}}$. Let $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{k})>0$. From Prop. 23, we can suppose that $\nu_{k}>\inf_{n\in\mathbb{N}}\nu_{n}$ and so there exists $\delta>0$ such that $(\nu_{k}-\delta,\nu_{k})\subset I_{X^{\mathrm{in}},T}^{\varphi}$. As $\alpha$ is continuous and strictly increasing on $I_{X^{\mathrm{in}},T}^{\varphi}$, for $\varepsilon:=\alpha(\nu_{k})/2$, there exists $\eta\in(0,\delta]$, such that $0<\nu_{k}-s\leq\eta$ implies that $\varepsilon<\alpha(s)<\alpha(\nu_{k})+\varepsilon$. Now let us consider $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{j}(x))>\nu_{k}-\eta$. We have either $\nu_{k}>\varphi(T^{j}(x))$ or $\varphi(T^{j}(x))\geq\nu_{k}$. In the first case, we have, by hypothesis, $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))>\varepsilon$. In the second case, as $\alpha$ is increasing and $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))\geq\alpha(\nu_{k})>\varepsilon$.
@@
Now, assume that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at some $k\in{\mathrm{R}}_{\nu}^{s}$. Then, from Prop.23, $g$ is also $(X^{\mathrm{in}},T,\varphi)$-compatible at $k=\mathbf{K}_{\nu}$. Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{\rm opt}-\eta$, we have $g(T^{j}(x))>\varepsilon$. Then we can define : $\alpha:s\mapsto s-\nu_{\rm opt}+\min\{\eta,\varepsilon\}$. Let $x\in X^{\mathrm{in}}$ and $j\in\mathbb{N}$. Then either $\varphi(T^{j}(x))\leq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$ or $\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}$. For the first case, $\alpha(\varphi(T^{j}(x)))\leq 0\leq g(T^{j}(x)))$. For the second, $0<\alpha(\varphi(T^{j}(x)))\leq\min\{\eta,\varepsilon\}$. By assumption on $g$, $g(T^{j}(x))>\varepsilon$. Note that the set $\{(x,j)\in X^{\mathrm{in}}\times\mathbb{N}:\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty. Finally, as $\alpha$ is strictly increasing, for all $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$, we have $\alpha(\nu_{k})>0$. Note that the set $\{k\in\mathbb{N}:\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty as $[\mathbf{K}_{\nu},\mathbf{K}^{s}_{\nu}]\cap\mathbb{N}$ is included in it. ∎
Example 9.
Let consider the Opt-Lyapunov function $V$ of Example 5. We prove that $V$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
We define for all $s\in\mathbb{R}$, $\alpha(s):=s-p^{2}/3+\min\{\zeta,\varepsilon\}$ where $0<\zeta<p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ and $\varepsilon:=\min\{\varepsilon_{k},k\in\operatorname{Argmax}{\nu}\}$ where $\varepsilon_{k}=\inf\{V(T^{k}(x)):x\in X^{\mathrm{in}}\}$. The scalar $\varepsilon$ is strictly positive as the function $[\mu,1/2]\ni y\mapsto V(T^{k}((2,1)y)$ is continuous and strictly positive for all $k\in\mathbb{N}$ and $\operatorname{Argmax}(\nu)$ is a finite set of integers. The function $\alpha$ is clearly strictly increasing and continuous on $\mathbb{R}$.
We have for all $j\in\mathbb{N}$ and for all $x\in X^{\mathrm{in}}$, $\alpha(\varphi(T^{j}(x)))=\varphi(T^{j}(x))-p^{2}/3+\min\{\zeta,\varepsilon\}<\varphi(T^{j}(x))-p^{2}/3+p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$. Then $\alpha(\varphi(T^{j}(x)))$ is strictly negative for all $j\notin\operatorname{Argmax}(\nu)$. Now let $j\in\operatorname{Argmax}(\nu)$. Suppose that $\varphi(T^{j}(x))>p^{2}/3-\min\{\zeta,\varepsilon\}$. As we have $\varphi(T^{j}(x))\leq p^{2}/3$, we get $\alpha(\varphi(T^{j}(x)))\leq\varepsilon\leq V(T^{k}(x))$ for all $k\in\operatorname{Argmax}(\nu)$ and all $x\in X^{\mathrm{in}}$. This leads to $\alpha(\varphi(T^{j}(x)))\leq V(T^{j}(x))$. On the other hand, $\alpha(\varphi(T^{j}(x)))\leq 0$ for all $j\in\operatorname{Argmax}(\nu)$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))\leq p^{2}/3-\min\{\eta,\varepsilon\}$. We have proved that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$. Moreover, as the set of $\{x\in X^{\mathrm{in}}:\varphi(T^{j}(x))>p^{2}/3-\min\{\eta,\varepsilon\}\}$ is nonempty for all $j\in\operatorname{Argmax}(\nu)$, we conclude that for all $j\in\operatorname{Argmax}(\nu)$, there exists $x\in X^{\mathrm{in}}$ such that $0<\alpha(\varphi(T^{j}(x)))$ and as $\alpha$ is increasing $0<\alpha(\nu_{j})$ for all $j\in\operatorname{Argmax}(\nu)$ and $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$ making $V$ $(X^{\mathrm{in}},T,\varphi)$-compatible.
Actually, by simple computations, we have $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$. One can compute $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ as follows. Let $n\in\mathbb{N}$. We have, for all $y\in\mathbb{R}_{+}^{*}$, $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$. Thus $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$ is strictly negative if and only if $n\geq n_{y}^{*}:=\max\{\lfloor\ln(4p/(15y))/\ln(3/2)\rfloor+1,0\}$. Then for all $y\in\mathbb{R}_{+}^{*}$, $\mathbb{N}\ni n\mapsto\varphi_{n}(y)$ is maximized at $n_{y}^{*}$. We deduce that $n_{1/2}^{*}=\lfloor\ln(8p/15)/\ln(3/2)\rfloor+1$ and, as $\mu\in(0,1/3)$, $n_{\mu}^{*}=\lfloor\ln(4p/(15\mu))/\ln(3/2)\rfloor+1$. This implies that $n_{1/2}^{*}\leq\underline{n}-1$ and thus we have $\nu_{n}\leq\nu_{n_{1/2}^{*}}$ for all $n\leq\underline{n}-1$. This also implies that $n_{\mu}^{*}<\overline{n}$ and so $\nu_{n}<\nu_{\overline{n}}$ for all $n>\overline{n}$. Finally, $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}=\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\}$.
5.2.4 Direct and Converse Opt-Lyapunov Theorems
@@
We present both direct and converse Opt-Lyapunov theorems using the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion. More precisely, the direct version aims to prove that if a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function exists (and such that $\overline{N_{T}}(V)>0$), then we can construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$. The converse version exhibits a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function when a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$. Moreover, we are able to produce a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for this Opt-Lyapunov function.
@@
We remark that from Prop. 14, if $V$ is an Opt-Lyapunov for $(X^{\mathrm{in}},T)$ then $\{T^{k}(x):k\in\mathbb{N},x\in X^{\mathrm{in}}\}\subseteq\operatorname{dom}(V)$.
@@
If $N_{V}(T)=0$ then for all $x\in\mathbb{R}^{d}$ such that $V(x)>0$, $V(T(x))=0$. This also implies that $V(T^{k}(x))=0$ for all $k\in\mathbb{N}^{*}$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. Thus, for all $k\in\mathbb{N}^{*}$, $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))=0$ and as $\alpha$ is increasing and continuous, $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}^{*}$. As for some $j\in{\mathrm{R}}_{\nu}^{s}$, $\alpha(\nu_{j})>0$, this leads to $j=0$ and as $\alpha$ is strictly increasing and continuous on $I_{X^{\mathrm{in}},T}^{\varphi}$ then for all $k\in\mathbb{N}^{*}$, $\nu_{0}>\nu_{k}$ and $\mathbf{K}^{s}_{\nu}=0$. This also implies that $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$. ∎
@@
Assume that there exists a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov $V$ such that $N_{V}(T)>0$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$. Let $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)$ where $\alpha(I)=[0,1]$. We define the function $h:x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$. Then $(h,N_{V}(T))\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$.

### v4

@@ 5.1.2 Operator Ratios From Proper Positive Semidefinite Functions
We say that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if there exists some $k\in{\mathrm{R}}_{\nu}^{s}$ at which $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
Example 8 ($V_{\ell}$ in (32) is $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-compatible).
We have to prove that there exists $k\in{\mathrm{R}}_{\omega}^{s}$, $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$, for all $j\in\mathbb{N}$ such that $\varphi_{\ell}(L^{j}x)>\omega_{k}-\eta$, we have $V_{\ell}(L^{j}x)>\varepsilon$.
Recall that $\overline{\lim}_{n}\omega_{n}=\lambda_{2}/c$ and thus $k\in{\mathrm{R}}_{\omega}^{s}$ iff $\omega_{k}>\lambda_{2}/c$. Let $k\in{\mathrm{R}}_{\omega}^{s}$ and let $\eta>0$ any real such that $\omega_{k}-\eta>\lambda_{2}/c$. Let $K\in\mathbb{N}$ such that $\omega_{j}\leq\lambda_{2}/c$ for all $j>K$. Let $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}_{\ell}$ such that $\varphi_{\ell}(L^{j}(x))>\omega_{k}-\eta$. It follows that $j\in\{0,\ldots,K\}$. From the discussion in Subsection 3.3.1, $\pi_{2}(L^{j}x)\leq\max\{\pi_{2}(x),\pi_{2}(L^{K}x)\}$. Moreover, as for all $x\in\mathbb{R}^{2}$ and all $k\in\mathbb{N}$, $\pi_{2}(L^{k}x)=(\lambda_{2}-\lambda_{1})^{-1}(\lambda_{1}^{k}\mathfrak{a}(x)+\lambda_{2}^{k}\mathfrak{b}(x))$, it is easy to see that $\pi_{2}(L^{j}x)\leq\max\{u_{2},\pi_{2}(L^{K}x^{uu})\}$. We conclude that, for all $x\in X^{\mathrm{in}}$ and all $j\in\mathbb{N}$ such that $\varphi_{\ell}(L^{j}x)>\omega_{k}-\eta$, we have
$\begin{array}[]{ll}V_{\ell}(L^{j}x)=\dfrac{C}{(\pi_{1}(L^{j}x))^{2}+(\pi_{2}(L^{j}x))^{2}}&=\dfrac{C}{(\pi_{2}(L^{j}x))^{2}((\varphi_{\ell}(L^{j}x))^{2}+1)}\\ &\geq\dfrac{C}{(\max\{u_{2},\pi_{2}(L^{K}x^{uu}))\})^{2}(\omega_{\rm opt}^{2}+1)}:=\varepsilon>0\end{array}$
This proves that $V_{\ell}$ is an Opt-Lyapunov function $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-compatible. ∎
@@
Suppose that there exists $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that $\alpha(\nu_{k})>0$ for some $k\in{\mathrm{R}}_{\nu}^{s}$ and $\alpha(T^{k}(x))\leq g(T^{k}(x))$ for all $k\in\mathbb{N}$ all $x\in X^{\mathrm{in}}$. Let $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{k})>0$. From Prop. 11, we can suppose that $\nu_{k}>\inf_{n\in\mathbb{N}}\nu_{n}$ and thus there exists $\zeta>0$ such that $(\nu_{k}-\zeta,\nu_{k})\subset I_{X^{\mathrm{in}},T}^{\varphi}$. As $\alpha$ is continuous and strictly increasing on $I_{X^{\mathrm{in}},T}^{\varphi}$, for $\varepsilon:=\alpha(\nu_{k})/2$, there exists $\eta\in(0,\zeta]$, such that $0<\nu_{k}-s\leq\eta$ implies that $\varepsilon<\alpha(s)<\alpha(\nu_{k})+\varepsilon$. Now let us consider $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{j}(x))>\nu_{k}-\eta$. We have either $\nu_{k}>\varphi(T^{j}(x))$ or $\varphi(T^{j}(x))\geq\nu_{k}$. In the first case, we have, by hypothesis, $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))>\varepsilon$. In the second case, as $\alpha$ is increasing and $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))\geq\alpha(\nu_{k})>\varepsilon$.
@@
Now, assume that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at some $k\in{\mathrm{R}}_{\nu}^{s}$. Then, from Prop.11, $g$ is also $(X^{\mathrm{in}},T,\varphi)$-compatible at $k=\mathbf{K}_{\nu}$. Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{\rm opt}-\eta$, we have $g(T^{j}(x))>\varepsilon$. Then we can define: $\alpha:s\mapsto s-\nu_{\rm opt}+\min\{\eta,\varepsilon\}$. Let $x\in X^{\mathrm{in}}$ and $j\in\mathbb{N}$. Then either $\varphi(T^{j}(x))\leq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$ or $\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}$. For the first case, $\alpha(\varphi(T^{j}(x)))\leq 0\leq g(T^{j}(x)))$. For the second, $0<\alpha(\varphi(T^{j}(x)))\leq\min\{\eta,\varepsilon\}$. By assumption on $g$, $g(T^{j}(x))>\varepsilon$. Note that the set $\{(x,j)\in X^{\mathrm{in}}\times\mathbb{N}:\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty. Finally, as $\alpha$ is strictly increasing, for all $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$, we have $\alpha(\nu_{k})>0$. Note that the set $\{k\in\mathbb{N}:\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty as $\mathbf{K}_{\nu},\mathbf{K}^{s}_{\nu}]\cap\mathbb{N}$ is included in it. ∎
Example 9 (A $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$).
From Th 4 and Example 8, there exists a $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$. We will see in Theorem 5 that this certificate is required to construct an element $(h,\beta)\in\mathbb{\Gamma}(\omega)$ from an Opt-Lyapunov function.
Thus, we construct a certificate of compatibility for $V_{\ell}$ proposed in (32). From the definition of $V_{\ell}$, we can easily establish a link between $V_{\ell}$ and $\varphi_{\ell}$. Indeed, for all $x\in(\mathbb{R}_{+}^{*})^{2}$:
$V_{\ell}(x)=\dfrac{C}{x_{1}^{2}+x_{2}^{2}}=\dfrac{C}{x_{1}^{2}(1+(\varphi_{\ell}(x))^{-2})}.$
The functions of the form $s\mapsto C/(M^{2}(1+s^{-2}))$ are strictly increasing and continuous. One way to use these functions as lower bounds of $V_{\ell}$ is to find an upper bound of the set $\{\pi_{1}(L^{k}x):x\in X^{\mathrm{in}}_{\ell},k\in\mathbb{N}\}$. However, from the discussion in 3.3.1, $\{\pi_{1}(L^{k}x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}$ is unbounded. Moreover, of the form $s\mapsto C/(M^{2}(1+s^{-2}))$ are strictly positive, whereas a certificate of compatibility must take negative values for all $k\in\mathbb{N}$ such that $\omega_{k}<\overline{\lim_{n}}\omega_{n}$.
Then, the certificate is defined as a piecewise function on $\mathbb{R}$. The pieces are one interval and its complement. The interval has a lower bound that is strictly greater than the limit superior. Hence, the certificate takes the form $s\mapsto C/(M^{2}(1+s^{-2}))$ on the interval. The definition of the interval permits us to consider a bounded subset of $\{\pi_{1}(L^{k}x):(x,k)\in X^{\mathrm{in}}_{\ell}\times\mathbb{N}\}$ and thus compute $M$. On the complement of the interval, our certificate is an affine form that is proved to be a lower bound of $V_{\ell}$.
First, from the analysis of $\mathfrak{a}$, we have either $\omega_{0}\in{\mathrm{R}}_{\omega}^{s}$ or $\omega_{1}\in{\mathrm{R}}_{\omega}^{s}$. Then, we define
$z:=\max\{\omega_{0},\omega_{1}\}.$
[…]
$N(x):=\left\lfloor\dfrac{\ln\left(\dfrac{\mathfrak{b}(x)(c(z-\eta)-\lambda_{2})}{|\mathfrak{a}(x)|(c(z-\eta)-\lambda_{1})}\right)}{\ln(\left|\lambda_{1}/\lambda_{2}\right|)}\right\rfloor$
For a given $x\in X_{\ell}$, the integer $N(x)$, if nonnegative, determines an upper bound of the maximal integer $k$ such that $\varphi_{\ell}(L^{k}x)\geq z-\eta$. Hence, the integer $\overline{N}$ defined as follows
[…]
is an upper bound for $K$. Moreover, it can be proved that $\overline{N}$ is nonnegative. Then, we can compute $M=\max_{x\in X^{\mathrm{in}}_{\ell}}\pi_{1}(L^{\overline{N}}x)=\pi_{1}(L^{\overline{N}}x^{uu})$. Actually, $\overline{N}$ is an upper bound of the greatest maximizer $\mathbf{K}_{\omega}$. However, we continue as the goal is to produce a $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-certificate of compatibility. We are able to define:
$\alpha_{1}:s\mapsto\dfrac{C}{M^{2}(1+s^{-2})}.$
Now let us take any $\gamma$ greater than
$\max\left\{\max_{s\in[z-\eta,z]}\alpha_{1}^{\prime}(s),\dfrac{V_{\ell}(x^{*})}{\eta}\right\}$
where $x^{*}\in X^{\mathrm{in}}_{\ell}\cup L(X^{\mathrm{in}}_{\ell})$ is the vector such that $z=\varphi_{\ell}(x^{*})$. Note that as $\alpha_{1}$ is strictly increasing and $V_{\ell}$ is strictly positive, $\gamma$ is strictly positive. It follows that $\alpha_{1}(s)\geq\gamma(s-z)+\alpha_{1}(z)$. Moreover, there exists $u\in[z-\eta,z)$ such that $\gamma(u-z)+\alpha_{1}(z)=0$. We conclude that
$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{C}{M(1+s^{-2})}&\text{ if }s\geq z\\ \gamma(s-z)+\dfrac{C}{M^{2}(1+z^{-2})}&\text{ if }s<z\end{array}\right$.
is a $(X_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$. As for all $s\in[z-\eta,z]$, we have
$\alpha_{1}^{{}^{\prime}}(s)=\dfrac{2Cs^{-3}}{(M^{2}(1+s^{-2}))^{2}}\leq\dfrac{2C(z-\eta)^{-3}}{(M^{2}(1+z^{-2})},$
a candidate $\gamma$ could be
$\max\left\{\dfrac{2C(z-\eta)^{-3}}{(M^{2}(1+z^{-2}))^{2}},\dfrac{V_{\ell}(x^{*})}{\eta}\right\}.$
Applying those computations to our numerical data of Subsection 3.4, we have $z=\max\{\omega_{0},\omega_{1}\}=42$ and
$\eta=0.5(42-4)=19$
We have $\mathfrak{b}(x^{lu})=55$, $\mathfrak{a}(x^{lu})=95$, $\mathfrak{b}(x^{ul})=25$ and $\mathfrak{a}(x^{ul})=-10$. Hence,
$N(x^{lu})=\left\lfloor\dfrac{\ln\left(\dfrac{319}{665}\right)}{-\ln(2)}\right\rfloor=1\text{ and }N(x^{ul})=\left\lfloor\dfrac{\ln\left(\dfrac{19}{10}\right)}{-\ln(2)}\right\rfloor=-1$
Thus, the integer $\overline{N}$ is 1. We remark that the integer $N(x^{ul})=-1$ which means that to have $\varphi_{\ell}(L^{k}x^{ul})\geq 42-19=23$, $k$ has to be negative. Indeed, for all $k\in\mathbb{N}$, $\varphi_{\ell}(L^{k}x^{ul})<23$. We could stop here as $\overline{N}$ is an upper bound of the greatest integer maximizer of ($DPCP_{\ell}$).
We have to compute $M=\pi_{1}(Lx^{uu})^{2}=240^{2}=57600$. We have $x^{*}=Lx^{lu}=(210,5)^{\intercal}$. Then
$\dfrac{V_{\ell}(x^{*})}{\eta}=\dfrac{125}{(19(210^{2}+5^{2}))}=\dfrac{1}{6707}\text{ and }\dfrac{2C(z-\eta)^{-3}}{M(1+z^{-2})^{2}}=\dfrac{250\times 23^{-3}}{57600(1+42^{-2})^{2}}=\dfrac{21609}{60644708120}$
and thus $\gamma=1/6707$.
We conclude that
$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{125}{57600(1+s^{-2})}&\text{ if }s\geq 42\\ \\ \dfrac{(s-42)}{6707}+\dfrac{49}{22592}&\text{ if }s<42\end{array}\right$.
5.3 Direct and Converse Opt-Lyapunov Theorems
@@
We present both direct and converse Opt-Lyapunov theorems using the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion. More precisely, the direct version aims to prove that if a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function exists (and such that $\overline{N_{T}}(V)>0$), then we can construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ that is useful for $\nu$. The converse version exhibits a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function when a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$. Moreover, we can produce a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for this Opt-Lyapunov function.
@@
We note that from Prop. 2, if $V$ is an Opt-Lyapunov for $(X^{\mathrm{in}},T)$ then $\{T^{k}(x):k\in\mathbb{N},x\in X^{\mathrm{in}}\}\subseteq\operatorname{dom}(V)$.
@@
If $N_{V}(T)=0$ then for all $x\in\mathbb{R}^{d}$ such that $V(x)>0$, $V(T(x))=0$. This also implies that $V(T^{k}(x))=0$ for all $k\in\mathbb{N}^{*}$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. Thus, for all $k\in\mathbb{N}^{*}$, $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))=0$ and as $\alpha$ is increasing and continuous, $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}^{*}$. For some $j\in{\mathrm{R}}_{\nu}^{s}$, $\alpha(\nu_{j})>0$, this leads to $j=0$ and as $\alpha$ is strictly increasing and continuous on $I_{X^{\mathrm{in}},T}^{\varphi}$ then for all $k\in\mathbb{N}^{*}$, $\nu_{0}>\nu_{k}$ and $\mathbf{K}^{s}_{\nu}=0$. This also implies that $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$. ∎
@@
Assume that there exists a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov $V$ such that $N_{V}(T)>0$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$. Let $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)$ where $\alpha(I)=[0,1]$. We define the function
$h:[0,1]\ni x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right).$ (35)
Then $(h,N_{V}(T))\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$.